I am having problem answering these questions. Please kindly help me.
1: For what real number k, does 3x2 – k factor as 3(x – 2)(x + 2)?

2: Which of the following products is not equivalent to the others? Explain your answer.
1. (2x – 4)(x + 3)
2. 2(x – 2)(x + 3)
3. (x – 2)(2x + 6)
4. (2x – 4)(2x + 6)

Answer by stanbon(48551)

1: For what real number k, does 3x^2 – k factor as 3(x – 2)(x + 2)?
If 3x^2-k = 3(x-2)(x+2)
then 3x^2-k = 3(x^2-4)
and 3x^2-k = 3x^2-12
Therefore k would have to equal 12
-------------------------------------------
2: Which of the following products is not equivalent to the others? Explain your answer.
1. (2x – 4)(x + 3)= 2x^2+2x-12
2. 2(x – 2)(x + 3)= 2(x^2+x-6)=2x^2+2x-12
3. (x – 2)(2x + 6)= 2x^2+2x-12
4. (2x – 4)(2x + 6)= 4x^2+4x-24
Ans: #4 is double each of the others
=============
Cheers,
Stan H.